OpenStudy (anonymous):

The Jeffersons are hiring a landscaper. The plants they buy will cost $150. Labor costs $40 per hour. The Jeffersons can spend no more than $450 on the project. How many hours of labor can the Jeffersons afford? Let h represent the hours of labor. no more than 6 hours 8 hours or less at most 12 hours a maximum of 15 hours

OpenStudy (anonymous):

40h = 450-150 40 h = 300 h=300/40= 7.5 hours of labor

OpenStudy (anonymous):

Thanks!